On October 1, 2010, Hess Company places a new asset into service. The cost of the asset is $40,000 with an estimated 5-year life and $10,000 salvage value at the end of its useful life. What is the book value of the plant asset on the December 31, 2010, balance sheet assuming that Hess Company uses the double-declining-balance method of depreciation?

$26,000.

$30,000.

$36,000.

$38,000.
Answer #1

Rate of double decline depreciation = 5 years = 40%

Depreciation for 2010 = 40000*40%*3/12 = $4000

Book value on the December 31, 2010 = 40000-4000 = $36000